OpenStudy (itsbribro):

I really don't understand. PLEASE HELP ^-^ Which is an equation in point-slope form for the given point and slope? Point: (2, -6) Slope: -3/4

OpenStudy (itsbribro):

A. \[y-6=-\frac{ 3 }{ 4 }(x+2)\] B. \[y-6=\frac{ 3 }{ 4 }(x-2)\] C. \[y+6=-\frac{ 3 }{ 4 }(x-2)\] D. \[y+6=-\frac{ 3 }{ 4 }(x+2)\]

OpenStudy (catlover5925):

\[y - y _{1}=M(x-x _{1)}\]

OpenStudy (catlover5925):